OpenStudy (anonymous):

the difference between the digits in a two-digit numeral is 3. The units digit is twice the tens digit. Find the numeral.

OpenStudy (anonymous):

we could write the first words in math as \(x-y=3\) i.e. the difference is 3

OpenStudy (anonymous):

then the second one means \(y=2x\)

OpenStudy (anonymous):

oh that isn't going to work, lets say \(y-x=3\) and \(y=2x\)

OpenStudy (anonymous):

that makes \(2x-x=3\) or \(x=3\) and so \(y=6\) making your number \(36\)

OpenStudy (anonymous):

Let the tens digit be x. Difference is 3. Units digit is twice the tens digit so.. x - 2x = 3 -x = 3 x = -3 Tens digit is -3 and one digit -6. Number is -36.

OpenStudy (anonymous):

How do ypu the equation?

OpenStudy (anonymous):

You*